OpenStudy (anonymous):

What's wrong with\[\int_{-1}^1\frac{dx}{x}=\ln|1|-\ln|-1|=0\]?

OpenStudy (anonymous):

Yes, but I'm told that this is incorrect.

OpenStudy (callisto):

I guess you have to 'separate' the integral into 2 parts (from -1->0 and 0->1) and use lim to find the area?! It seems i've encountered some problem like this and you have to use limit to do it..

OpenStudy (anonymous):

Gotcha.

OpenStudy (anonymous):

Because you're integrating \(\ln x\) in a region underfined for the reals.

OpenStudy (callisto):

the term is something like 'improper integral'.. I'm not sure and i couldn't find that post.. sorry :(

OpenStudy (callisto):

Hmm the post is this http://openstudy.com/study#/updates/4f6ac4a6e4b014cf77c7e6e1 That's something similar to your problem

myininaya (myininaya):

Yes @callisto 's is right You have to break this up

myininaya (myininaya):

If both of the parts converge then find the sum of what both converge to and that is your answer

myininaya (myininaya):

If one part diverges then the whole thing diverges

myininaya (myininaya):

\[\lim_{a \rightarrow 0^-} \int\limits_{-1}^{a} \frac{1}{x} dx+\lim_{b \rightarrow 0^+} \int\limits_{b}^{1} \frac{1}{x} dx\]

myininaya (myininaya):